Calculate the required rate of return for Mudd Enterprises assuming that investors expect a 5.0% rate of inflation in the future. The real risk-free rate is 1.0%, and the market risk premium is 7.0%. Mudd has a beta of 1.8, and its realized rate of return has averaged 8.5% over the past 5 years. Round your answer to two decimal places.
Answer #1

Risk Free Rate = 0.05 + 0.01 = 6.00%

Using CAPM Model,

Rate of return = 0.06 + 1.80(0.07)

Rate of Return = 18.60%